### v1

@@ 3 Catastrophe Bond Data
4 Empirical Analysis
In this section, we provide empirical analyses of CAT bond pricing. Sections 4.1 and 4.2 present results obtained using the linear regression model and two ensemble learning methods, respectively.
4.1 Linear Regression Models
Our analysis begins with constructing an extensive linear regression model, incorporating a range of predictors considered in existing studies (Braun,, 2016; Gürtler et al.,, 2016; Beer and Braun,, 2022). The list of predictors and their descriptions are provided in Table B.1. The linear regression model takes the following form:
@@ 4.1 Linear Regression Models
To address the potential overfitting problem, we employ variable selection techniques, including LASSO and forward/backward stepwise selection. Table 4.1 presents the estimation results for the full model (4.1), LASSO regression, and the optimal model (with the smallest AIC value) resulting from the stepwise selection. All models are applied to the full sample of 765 bonds. In the full model, the $p$-value of the coefficient for EL is close to zero, indicating its significance in predicting CAT bond spreads. The coefficient of EL is 179, indicating that a 1% increase in EL would lead to a 1.79% increment of the bond spread, all else equal. The identified strong correlation between EL and bond spread is consistent with findings in the literature. Nevertheless, approximately one-third of the parameters are statistically insignificant, indicating the need for variable selection. We also find that the coefficient estimate for PFL is negative and not statistically significant, which seems counter-intuitive. A positive correlation between PFL and CAT bond spreads is well-documented within the literature, as evidenced by Lane, (2000), Wills and Sherris, (2010), and Beer and Braun, (2022). Our finding can be attributed to a high degree of collinearity between PFL and EL, as indicated by Figures 1(a) and 1(b). Such collinearity can lead to coefficient estimates that are unstable and also result in inflated standard errors. While collinearity may not compromise the overall predictive capacity, it poses challenges to the interpretability of the regression coefficients within the model. For the same reason, coefficients for the Guy Carpenter indices and their $p$-values should be interpreted with caution due to the multi-collinearity among them.
@@
The model derived from LASSO regression removes six variables, PFL, WIND, US, EU, GC.UK, and JP×$\times$×GC.AP, from the full model, resulting in slightly lower $R^{2}$ and adjusted $R^{2}$ compared to the full model. The optimal model from step-wise variable selection excludes PFL, SIZE, INDEM, WIND, US, EU, and JP, six of which were found to be insignificant in the full model. By excluding these non-significant variables, the optimal step-wise model achieves the highest adjusted $R^{2}$ among the three models assessed. Moreover, all variables included in the optimal model are statistically significant.
@@
The coefficients in the selected model exhibit the expected signs. For example, bonds covering earthquakes have on average lower spreads, as do bonds issued by Swiss Re or those with an investment-grade rating. It is worth noting that the negative coefficient associated with SR does not contradict the descriptive statistics in Table 3.2, where bonds issued by Swiss Re are shown to have higher spreads on average. This discrepancy arises because bonds issued by Swiss Re typically have a higher average EL, leading to higher spreads. However, when all else is equal, the linear regression model predicts that having Swiss Re as the issuer tends to lower bond spreads.
@@
Table 4.1: Estimation results of the linear regression model, including the full model (4.1), LASSO, and the optimal model selected using step-wise backward and forward variable selection, respectively. Significance levels: ‘***’ $<0.001$, ‘**’ $<0.01$, ‘*’ $<0.05$, ‘.’ $<0.1$.
Full Model LASSO Step-wise Selection
Coeff. Std. p-Val. Sig. Coeff. Coeff. Std. p-Val. Sig.
(Intercept) -149.58 120.39 0.21 -177.64 -174.13 104.46 0.10 .
EL 179.05 17.22 <2e-16 *** 165.41 165.68 4.23 <2e-16 ***
PFL -9.03 12.40 0.47 .
SIZE 0.11 0.09 0.21 0.05
TERM -3.34 0.93 0.00 *** -3.30 -3.12 0.88 0.00 ***
INDEM 14.62 23.15 0.53 1.19
WIND -1.64 28.51 0.95 .
EQ -182.26 31.86 0.00 *** -171.87 -176.69 28.16 0.00 ***
US -61.24 90.76 0.50 .
EU 20.64 198.08 0.92 .
JP -516.14 270.93 0.06 . -42.72
SR -65.96 23.34 0.00 ** -58.17 -68.53 22.91 0.00 **
IG -200.22 53.18 0.00 *** -187.51 -202.21 52.81 0.00 ***
ROLX 3.22 0.83 0.00 *** 3.34 3.23 0.82 0.00 ***
BBSPR 0.32 0.08 0.00 *** 0.26 0.32 0.08 0.00 ***
GC.GLOB 2.41 0.68 0.00 *** 2.37 2.39 0.67 0.00 ***
GC.US -2.14 0.50 0.00 *** -2.15 -1.89 0.42 0.00 ***
GC.AP -2.78 0.59 0.00 *** -1.76 -2.65 0.58 0.00 ***
GC.EU 7.84 2.76 0.00 ** 2.86 7.75 2.72 0.00 **
GC.UK -7.08 4.10 0.08 . . -7.26 4.06 0.07 .
JP×$\times$×GC.AP 2.70 1.52 0.08 . . 2.76 1.51 0.07 .
US×$\times$×GC.US 0.83 0.50 0.10 . 0.46 0.51 0.12 0.00 ***
EU×$\times$×GC.EU -1.73 1.50 0.25 -1.51 -1.59 0.30 0.00 ***
$R$-squared 0.775 0.772 0.7741
$R$-squared Adj 0.7684 0.7635 0.7693
4.2 Machine Learning Methods
@@
We now proceed to the application of Random Forest and XGBoost models to CAT bond pricing. In these models, we incorporate all the variables in Equation (4.1), except interaction terms. The exclusion of interaction terms is deliberate, as machine learning models inherently capture interactions between variables. We keep both PFL and EL in the fitting of Random Forest and XGBoost models since tree-based methods are less sensitive to collinearity. However, high collinearity can still affect the interpretation of tree-based models, particularly when it comes to understanding the importance or impact of individual predictor variables. We will discuss this further in Section 4.3.2. To avoid overfitting and improve model performance, machine learning methods often require careful tuning of hyperparameters. This tuning process employs an exhaustive grid search of a predefined set of hyper-parameters to identify the combination that optimizes cross-validation prediction accuracy. Key hyperparameters for both models and their optimal values obtained through grid search are presented in Table B.2.

### v2

@@ 2 Catastrophe Bond Data
3 Methodology
This section introduces the statistical learning methods used in our analysis. Section 3.1 formulates the CAT bond pricing problem, while Sections 3.2 and 3.3 introduce the linear model and ensemble machine learning methods, respectively. Finally, Section 3.4 outlines the Conformal Prediction approach.
3.1 Problem Formulation
Let us denote the spread of a CAT bond at issuance as $Y$, and let $\bm{X}=(X_{1},X_{2},\ldots,X_{p})$ represent $p$ distinct risk factors. This vector $\bm{X}$ encompasses variables that convey information about bond characteristics, market conditions, and economic indicators. We assume there exists a relationship between $Y$ and $\bm{X}$, expressed as:
$Y=f(\bm{X})+\epsilon.$
Assume that we have observed a collection of $n$ distinct data points from primary CAT bond market transactions. We use $x_{ij}$ to represent the value of the $j$th risk factor from the $i$th observation, where $i=1,2,\ldots,n$ and $j=1,2,\ldots,p$. Correspondingly, $y_{i}$ denotes the value of the final spread from the $i$th transaction. We can represent the observed dataset as $D=\{(\bm{x}_{1},y_{1}),(\bm{x}_{2},y_{2}),\ldots,(\bm{x}_{n},y_{n})\}$, where $\bm{x}_{i}=(x_{i1},x_{i2},\ldots,x_{ip})^{T}$.
The estimation of $f$ involves exploring the hypothesis space $H$, which is the set of all feasible functions, to find a function, denoted as $\hat{f}$, that can capture the relationship between $\bm{X}$ and $y$ as accurately as possible. This process, commonly referred to as learning, is guided by minimizing a loss function that quantifies the difference between the true output $(y_{1},y_{2},\ldots,y_{n})$ from the observed data and the predicted outputs $\bigl{(}\hat{f}(\bm{x}_{1}),\hat{f}(\bm{x}_{2}),\ldots,\hat{f}(\bm{x}_{n})\bigr{)}$.
3.2 Linear Regression
$f(\bm{X})=\beta_{0}+\beta_{1}X_{1}+\beta_{2}X_{2}+\cdots+\beta_{p}X_{p}.$
3.3 Ensemble Learning
This subsection introduces XGBoost, an ensemble learning method used in our analysis. An overview of ensemble methods is provided in Section 3.3.1, while Section 3.3.2 explore the specifics of XGBoost.
3.3.1 Overview
Ensemble learning is a widely adopted machine learning technique that leverages the power of multiple models (base learners) to collectively address a problem, resulting in improved accuracy and robustness compared to individual models. Instead of relying on a single function $f$ for estimation, ensemble learning combines the outputs of multiple functions that approximate $f$. In this study, we employ XGBoost, a well-established ensemble learning methods.
Ensemble learning uses base learners, often capable of capturing nonlinearity, such as decision trees, to model complex relationships between explanatory variables and the variable of interest. Consequently, ensemble learning frequently outperforms linear regression in model goodness-of-fit as well as prediction accuracy. However, the amalgamation of multiple models can make the output less interpretable. To better understand the model’s behavior, various strategies, including partial dependence plots and feature importance measures, have been developed to illustrate the significance of individual explanatory variables in model predictions.
3.3.2 XGBoost
3.4 Conformal Prediction
3.4.1 Prediction Interval
$P\bigl{(}y_{n+1}\in C_{\alpha}(\bm{x}_{n+1})\bigr{)}\geq 1-\alpha.$ (3.1)
3.4.2 Overview of Conformal Prediction
$P(u_{n+1}\leq\hat{q}_{1-\alpha}\{u_{1},\ldots,u_{n}\})\geq 1-\alpha,$
where $u_{(1)}\leq\ldots\leq u_{(n)}$ denote the order statistics of $u_{1},\ldots,u_{n}$, and $\lceil(n+1)(1-\alpha)\rceil$ is the smallest integer greater than or equal to $(n+1)(1-\alpha)$.
3.4.3 Jackknife+ Conformal Prediction Interval
  1. 1.
  2. 2.
  3. 3.
    Repeat Steps 1 and 2 for all $i=1,\ldots,n$.
  4. 4.
  5. 5.
  6. 6.
While the Jackknife method often empirically achieves $1-\alpha$ coverage, it does not guarantee finite sample coverage validity, indicating that, for a finite sample size, the interval may not contain the true value $100(1-\alpha)$of the time. Asymptotic validity is only ensured when non-trivial conditions are satisfied by the trained model (Steinberger and Leeb,, 2016). The Jackknife prediction interval may exhibit poor coverage when the trained model $\hat{f}$ is unstable due to its lack of finite sample coverage validity.
4 Linear Regression: Model Estimation and Implications
Our analysis begins with constructing an extensive linear regression model, incorporating a range of risk factors considered in existing studies (Braun,, 2016; Gürtler et al.,, 2016; Beer and Braun,, 2022). The list of risk factors and their descriptions are provided in Table B.1. The linear regression model takes the following form:
@@ 4 Linear Regression: Model Estimation and Implications
To address the potential overfitting problem, we employ variable selection techniques, including LASSO and forward/backward stepwise selection. Table 4.1 presents the estimation results for the full model (4.1), LASSO regression, and the optimal model (with the smallest AIC value) resulting from the stepwise selection. All models are applied to the full sample of 765 bonds.
In the full model, the $p$-value of the coefficient for EL is close to zero, indicating its strong explanatory power for CAT bond spreads. The coefficient of EL is 179, meaning that a 1% increase in EL would lead to a 1.79% increment of the bond spread, all else being equal. The strong correlation between EL and bond spread is consistent with findings in the literature. Nevertheless, approximately one-third of the parameters are statistically insignificant, indicating the need for variable selection.
@@
We also find that the coefficient estimate for PFL is negative and statistically insignificant, which appears counter-intuitive. A positive correlation between PFL and CAT bond spreads is well-documented within the literature, as evidenced by Lane, (2000), Wills and Sherris, (2010), and Beer and Braun, (2022). This result can be attributed to a high degree of collinearity between PFL and EL, as indicated by Figures 1(a) and 1(b). Such collinearity can lead to unstable coefficient estimates and inflated standard errors. While collinearity may not compromise the overall explanatory capacity of the model, it poses challenges for the interpretability of the regression coefficients. For the same reason, coefficients for the Guy Carpenter indices and their $p$-values should be interpreted with caution due to the multi-collinearity among them.
@@
The model derived from LASSO regression removes six variables – PFL, WIND, US, EU, GC.UK, and JP×$\times$×GC.AP – from the full model, resulting in slightly lower $R^{2}$ and adjusted $R^{2}$ compared to the full model. The optimal model from step-wise variable selection excludes PFL, SIZE, INDEM, WIND, US, EU, and JP, six of which were found to be insignificant in the full model. By excluding these non-significant variables, the optimal step-wise model achieves the highest adjusted $R^{2}$ among the three models assessed. Moreover, all variables included in the optimal model are statistically significant.
@@
The coefficients in the selected model exhibit the expected signs. For example, bonds covering earthquakes have on average lower spreads, as do bonds issued by Swiss Re or those with an investment-grade rating. It is worth noting that the negative coefficient associated with SR does not contradict the descriptive statistics in Table 2.2, where bonds issued by Swiss Re are shown to have higher spreads on average. This discrepancy arises because bonds issued by Swiss Re typically have a higher average EL, leading to higher spreads. However, when controlling for other factors, the linear regression model indicates that having Swiss Re as the issuer tends to lower bond spreads.
@@
Table 4.1: Estimation results of the linear regression model, including the full model (4.1), LASSO, and the optimal model selected using step-wise backward and forward variable selection, respectively. Significance levels: ‘***’ $<0.001$, ‘**’ $<0.01$, ‘*’ $<0.05$, ‘.’ $<0.1$. Full Model LASSO Step-wise Selection Coeff. Std. p-Val. Sig. Coeff. Coeff. Std. p-Val. Sig. (Intercept) -149.58 120.39 0.21 -177.64 -174.13 104.46 0.10 . EL 179.05 17.22 <2e-16 *** 165.41 165.68 4.23 <2e-16 *** PFL -9.03 12.40 0.47 . SIZE 0.11 0.09 0.21 0.05 TERM -3.34 0.93 0.00 *** -3.30 -3.12 0.88 0.00 *** INDEM 14.62 23.15 0.53 1.19 WIND -1.64 28.51 0.95 . EQ -182.26 31.86 0.00 *** -171.87 -176.69 28.16 0.00 *** US -61.24 90.76 0.50 . EU 20.64 198.08 0.92 . JP -516.14 270.93 0.06 . -42.72 SR -65.96 23.34 0.00 ** -58.17 -68.53 22.91 0.00 ** IG -200.22 53.18 0.00 *** -187.51 -202.21 52.81 0.00 *** ROLX 3.22 0.83 0.00 *** 3.34 3.23 0.82 0.00 *** BBSPR 0.32 0.08 0.00 *** 0.26 0.32 0.08 0.00 *** GC.GLOB 2.41 0.68 0.00 *** 2.37 2.39 0.67 0.00 *** GC.US -2.14 0.50 0.00 *** -2.15 -1.89 0.42 0.00 *** GC.AP -2.78 0.59 0.00 *** -1.76 -2.65 0.58 0.00 *** GC.EU 7.84 2.76 0.00 ** 2.86 7.75 2.72 0.00 ** GC.UK -7.08 4.10 0.08 . . -7.26 4.06 0.07 . JP×$\times$×GC.AP 2.70 1.52 0.08 . . 2.76 1.51 0.07 . US×$\times$×GC.US 0.83 0.50 0.10 . 0.46 0.51 0.12 0.00 *** EU×$\times$×GC.EU -1.73 1.50 0.25 -1.51 -1.59 0.30 0.00 *** $R$-squared 0.775 0.772 0.7741 $R$-squared Adj 0.7684 0.7635 0.7693
5 XGBoost: Model Estimation and Implications
@@
We now proceed to the application of XGBoost models to CAT bond pricing. In these models, we incorporate all the variables in Equation (4.1), except interaction terms. The exclusion of interaction terms is deliberate, as machine learning models inherently capture interactions between variables. We keep both PFL and EL in the fitting of XGBoost models since tree-based methods are less sensitive to collinearity. However, high collinearity can still affect the interpretation of tree-based models, particularly when it comes to understanding the importance or impact of individual predictor variables. We will discuss this further in Section 5.2. To avoid overfitting and improve model performance, machine learning methods often require careful tuning of hyperparameters. This tuning process employs an exhaustive grid search of a predefined set of hyper-parameters to identify the combination that optimizes cross-validation prediction accuracy. In our approach, a 5-fold cross-validation was used during the grid search to ensure robust hyperparameter selection. After identifying the optimal hyperparameters, the model was then trained on the entire dataset. Key hyperparameters for both models and their optimal values obtained through grid search are presented in Table B.2.
In the subsequent sections, we discuss the estimated XGBoost model and the implied impacts of various risk factors on CAT bond spreads. To achieve this, we employ feature importance and accumulated local effects (ALE) plots, both widely used tools for interpreting machine-learning-based models.
5.1 Relative Importance of Risk Factors
Figure 5.1 presents the feature importance plot which illustrates the relative importance of each risk factor in the trained XGBoost model, measured by its contribution to the model’s goodness of fit. The plot shows that expected loss is the most important variable in explaining CAT bond spreads, consistent with findings from existing studies and the linear regression models. Notably, unlike the linear regression models, PFL also plays an important role in improving the explanatory power of the XGBoost model, even when EL is included. This suggests that despite the high correlation between EL and PFL, the latter still provides additional valuable information for understanding CAT bond spreads. The linear regression model likely excludes PFL due to its nonlinear relationship with spreads, a characteristic that XGBoost is well-equipped to capture.
Figure 5.1: Feature importance plot for the XGBoost model.
Beyond EL and PFL, the roles of other variables in the model are also noteworthy. The Lane Financial LLC Synthetic RoL Index, $\mathrm{ROLX}$, ranks as the second most important variable in the XGBoost model. Additionally, the Guy Carpenter indices, $\mathrm{GC.EU}$ and $\mathrm{GC.GLOB}$, which track changes in reinsurance pricing for property catastrophe coverage in the European and global markets respectively, rank as the fifth and sixth most important features. These results align with the fact that CAT bond investors expect returns that reflect the risk they are assuming, often benchmarking these returns against the cost of traditional reinsurance. When reinsurance rates rise, CAT bond yields tend to increase as well. Moreover, reinsurance pricing reflects market perception of catastrophe risk. Changes in these prices can signal shifts in perceived risk, thereby influencing CAT bond pricing as investors and issuers reassess the risk-return profile. The importance of the three indices – ROLX, GC.EU, and GC.GLOB – is also evident in the linear regression model.
$\mathrm{SIZE}$ ranks as the fourth most important variable in the XGBoost model, despite being excluded from the stepwise-selected linear regression model (c.f. Table 4.1). The exclusion of $\mathrm{SIZE}$ from the linear regression model may stem from its nonlinear relationship with bond spreads and potential collinearity with other variables, which complicates isolating its unique contribution within a linear framework. The XGBoost model, however, can capture nonlinear relationships and interactions and is more robust to collinearity, as it selects features based on their contribution to improving performance across multiple decision trees. Consequently, the XGBoost model recognizes the importance of $\mathrm{SIZE}$ even when correlated risk factors are present. Finally, variables considered significant in the linear regression model, such as $\mathrm{IG}$ and $\mathrm{GC.US}$, contribute minimally to the goodness of fit in the XGBoost model.
5.2 Impacts of Individual Risk Factors
The feature importance plot provides an initial overview of each risk factor’s relative contribution. However, this plot alone does not reveal how each factor influences CAT bond spreads. To address this, we further examine the impact of the six most important factors on CAT bond spreads by analyzing their accumulated local effect (ALE) plots, as shown in Figure 5.2. All ALE plots are centered to have an average of zero for enhanced interpretability. ALE plots are similar to the widely used partial dependence plots (PDP) but correct for the bias that arises in PDP when explanatory variables are highly correlated (Apley and Zhu,, 2020). For instance, since EL and PFL are highly correlated, it is unlikely that a CAT bond would simultaneously have a very high EL and very low PFL, or vice versa. PDPs cannot adequately address such situations because they rely on the marginal distributions of the explanatory variables. Consequently, when evaluating the impact of EL on bond spreads, PDPs treat EL independently of other variables and may consider unrealistic cases where EL is high and PFL is very low. In contrast, ALE plots rely on conditional distributions, accounting for the correlation among variables, and leading to more accurate interpretations of the model. Therefore, ALE plots are a more appropriate choice for our analysis, considering the potential high correlations among risk factors.
From Figure 5.2, nonlinear patterns are observed in the ALE plots. First, EL shows the strongest impact on bond spreads. When EL is close to zero, the bond spread is approximately 350 bps lower than the average bond spread. As EL increases to 15%, the bond spread rises to approximately 1,250 bps above the average bond spread. For EL values below 5%, there is an almost linear relationship, where bond spreads increase with the expected loss. Beyond the 5% threshold, the relationship becomes less straightforward. A sharp increase in bond spread is observed as EL increases from 5% to 6.25%. However, as EL progresses from 6.25% to 10%, the increase in bond spread is minimal. For EL values above 10%, bond spreads experience a modest increase with EL. Caution is advised when interpreting the section with large EL values exceeding 10%, due to the limited number of observations in this range.
Second, the ALE plot for PFL indicates that higher probabilities of first loss are associated with increased bond spreads, consistent with existing literature (Beer and Braun,, 2022), which identifies a positive correlation between PFL and bond spread. However, our interpretation differs significantly. In Beer and Braun, (2022), where only PFL is used as an explanatory variable, the effect of EL is indirectly captured by PFL due to their high correlation. Consequently, the impact attributed to PFL in their model is a combined effect of both PFL and EL. In contrast, our XGBoost model incorporates both PFL and EL as explanatory variables, with EL being identified as the most significant factor. Given their strong correlation, much of the information represented by PFL is also encompassed within EL. As a result, PFL contributes to the XGBoost model with information that is not already accounted for by EL. Therefore, the XGBoost model effectively utilizes all the information contained in EL and PFL. The impact of this additional information in PFL is captured in the ALE plot.
The impact of PFL on the bond spreads is most significant at lower levels of PFL and diminishes to almost zero as PFL exceeds 5%. This indicates that at lower PFL values, investors are sensitive to changes in PFL and demand a higher spread for an increased PFL. However, as PFL reaches higher levels, further increases in PFL appear to only have a moderate impact on bond spread. A likely explanation for this observation is that when PFL is high, EL tends to be high as well due to their correlation. In such cases, investors may become less sensitive to small increases in PFL, shifting their focus more towards EL as the more influential factor in their decision-making process.
Third, the impacts of ROLX and GC.EU on bond spreads also exhibit clear non-linear patterns. The impact of GC.GLOB on the bond spread is similar to but less significant than that of ROLX and GC.EU. Both ROLX and GC.EU have a modest influence on bond spread when their values are below 120, indicating a relatively soft reinsurance market. In a soft market, there is generally an abundance of capital and a competitive environment. Investors might also perceive lower risk in this environment, resulting in a lower impact of changes in reinsurance market conditions on CAT bond spreads. Additionally, CAT bonds can be viewed as alternatives to traditional reinsurance products. Reinsurance remains the dominant risk transfer mechanism for catastrophe risks, and the choice between reinsurance and CAT bonds reflects a trade-off between their relative benefits and costs (Subramanian and Wang,, 2018). In a soft reinsurance market, where reinsurance capacity is ample and premiums are low, entities with hedging needs are more inclined to opt for reinsurance products rather than issuing CAT bonds. Small increases in reinsurance premiums are unlikely to push these entities toward the CAT bond market, as the increased premiums remain affordable. Therefore, CAT bond spreads remain low and relatively insensitive to reinsurance market conditions when the market is soft.
As the reinsurance market hardens, indicated by increased ROLX and GC indices, capital becomes tighter, demand for coverage rises, and the perception of risk intensifies. Investors seek higher returns for assuming the same risks, leading to a more pronounced impact of ROLX and GC.EU on bond spreads, as shown by the steep increase in their corresponding ALE plots. Due to limited capacity and high premiums in the reinsurance market, more entities with hedging demands turn to CAT bonds to transfer their risks. If reinsurance premiums continue to rise, more entities may shift towards the CAT bond market, driving up CAT bond spreads. This explains why CAT bond spreads are more sensitive to reinsurance market conditions when the market is hard.
Our analysis also indicates that CAT bond spreads tend to increase with issue size, particularly for issues smaller than 500 million US dollars. This observation contrasts the findings of Braun, (2016), who reported a negative impact of issue size on spreads. This difference results from the incorporation of GC indices in our model, which exhibit a moderate negative correlation (Pearson correlation of $-0.27$ for GC.EU and $-0.21$ for GC.GLOB) with the issue size. The negative correlation suggests that in a hardening market, characterized by tighter capital and higher perceived risks, CAT bond issue sizes tend to decrease. Therefore, issue size may serve as a weak indicator of market cycle conditions. In the absence of GC indices as predictors, as in the analysis by Braun, (2016), the observed negative impact of issue size on spreads could primarily reflect the effect of a hardening market. By incorporating GC indices as risk factors in our model, we can distinguish the specific impact of issue size on bond spreads from the effects captured by the GC indices. The ALE plot suggests that in a hard market, where issue size is often small, an increase in issue size is associated with an increase in bond spreads.
Overall, the ALE plots highlight significant nonlinear relationships between key risk factors and CAT bond spreads. Failing to account for these nonlinear patterns could potentially lead to inaccurate estimations of bond spreads.
5.3 Interaction Effects
Machine learning models like Random Forest and XGBoost naturally account for interaction effects between explanatory variables, providing an advantage over linear regression, where interactions must be explicitly defined. To visualize these interactions within the estimated XGBoost model, we utilize second-order ALE plots. While first-order ALE plots illustrate the main effect of individual risk factors, second-order ALE plots reveal the additional joint impact of pairs of risk factors on CAT bond spreads.
As an illustration, Figure 5.3 displays second-order ALE plots for four pairs of risk factors: EL & ROLX, EL & SIZE, EL & GC.EU, and EL & GC.GLOB. The contour values on these plots indicate the strength of the joint effects. All four plots demonstrate clear interaction effects between the respective pairs of risk factors, as the contour values vary significantly in each plot. The magnitude of these values is smaller compared to the ALE values depicted in Figure 5.2, suggesting that while these second-order effects are significant, they are generally smaller than the main effects of the individual risk factors.
(a) EL (x-axis) and ROLX (y-axis)
(b) EL (x-axis) and SIZE (y-axis)
(c) EL (x-axis) and GC.EU (y-axis)
(d) EL (x-axis) and GC.GLOB (y-axis)
Figure 5.3: Second-order ALE plots of four selected predictor pairs.
We first examine Figure 3(a), which displays the interaction between EL and ROLX. At an EL value of 2%, an increase in ROLX from 120 to 140 does not change the spread. In contrast, at an EL of 8%, the same increase in ROLX results in an 80 bps decrease in the spread, thereby suggesting a negative interaction between EL and ROLX. It is important to note that the second-order ALE effect is distinct from the main effect. To fully understand the impact of ROLX on the spread given the EL value, one must also consider the main effect of an increase in ROLX, which is positive within the range of 120–140. Specifically, an increase in ROLX from 120 to 140 results in approximately a 140 bps increase in the spread of a bond with a 2% EL, as illustrated in Figure 5.2. However, this increase translates to only a 60 bps (140 bps -- 80 bps) spread increase when EL is 8%. Therefore, for bonds with a higher EL, the impact of ROLX on the spread remains positive but weakened after combining the main and second-order effects.
Analysis of Figure 3(b) indicates that at an EL value of 2%, an increase in SIZE from 150 to 250 has a minimal effect on the spread. However, the same increase in SIZE results in approximately 45 bps decrease in the spread when EL is 10%, suggesting a negative interaction between EL and SIZE. We observed a positive main effect of SIZE on bond spread in Figure 5.2. The negative interaction between EL and SIZE suggests that an increase in issue size leads to a less pronounced rise, or even a decrease, in the spread for bonds with a higher EL. Given that the main effect of SIZE on the spread is relatively modest, as depicted in Figure 5.2, it is conceivable that the combined main and interaction effects of certain increases in SIZE could result in a negative impact on the spread for bonds with a high EL.
Figures 3(c) and 3(d) depict the interactions between EL and the GC indices, GC.EU and GC.GLOB, respectively. These interactions exhibit a complex pattern, neither consistently positive nor negative. For instance, at an EL value of 2%, increasing GC.EU from 140 to 160 does not affect the spread. At an EL of 4%, this same increase in GC.EU is associated with an increase in the spread. Interestingly, at an EL of 8%, an identical change in GC.EU results in a decrease in spread. The XGBoost model, with its tree-based structure, can capture these complex interactions, offering greater flexibility and higher explanatory power than models that assume linear relationships.
6 Prediction Performance and Conditional Impact
In this section, we examine the prediction performance of linear regression models and XGBoost models by analyzing their prediction accuracy and prediction intervals. In addition, we explore the conditional impacts of various factors by observing how changes in one factor influence the predictions and prediction intervals while fixing other factors. The prediction intervals are produced from the normally distributed error terms for the linear regression model and from employing the Jackknife+ prediction intervals introduced in Section 3.4.3 for XGBoost. Section 6.1 compares the prediction accuracy and prediction intervals generated by these two approaches. Section 6.2 further investigates the conditional impact of expected loss and its interactions with other factors.
6.1 Predictive Accuracy and Intervals
In statistical learning literature, it is common practice to randomly split the data into two parts: a training set used to train the model and a test set used to evaluate its predictive performance. However, relying on a single split can introduce randomness, making the analysis less robust. To address this, we perform a Monte Carlo cross validation involving 1,000 random data splits. For each split, 80% of the observations are randomly selected as the training set. This training set is first used for hyperparameter tuning through cross-validation, followed by training the final model using the identified optimal parameters. The remaining 20% of the data, which forms the test set, is then used strictly for prediction, allowing us to assess predictive performance. The mean squared errors (MSE), mean prediction interval length, and mean prediction interval coverage for the test set are calculated for each model and each split.